OpenStudy (lilg132):

lol whats the question

OpenStudy (anonymous):

question must be what question shall i ask

OpenStudy (anonymous):

hello joe!

OpenStudy (anonymous):

Um talk on chat don't talk here please

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{F_n}{F_{n-1}F_{n+1}}=?\] where F are fibonacci numbers <.< thats the question!

OpenStudy (anonymous):

yo satellite :)

OpenStudy (anonymous):

and what is the answer?

OpenStudy (anonymous):

1 i think...id have to think about it lol i answered it a few days ago.

OpenStudy (anonymous):

if you start the fib sequence at F_1=1, F_2=1, and F_n+1=F_n+F_n-1 i think thats rights.